OpenStudy (anonymous):

5^(2x+1)=8 Take the common logarithm of both sides of the equation. Use the power property to bring the exponent, (2x + 1), down as a factor. Show your work.

OpenStudy (anonymous):

\[\log_b(b^x) = x\] \[\log_5(5^{2x+1}) = \log_5(8)\] \[2x + 1 = \log_5(8)\] simplify for x

OpenStudy (wolf1728):

(2x +1) * log (5) = log (8)

OpenStudy (agent0smith):

@Euler271 good answer but you didn't follow the instructions.

OpenStudy (agent0smith):

Also, log base 5 of 8 isn't all that helpful now that you'll have to change the base :P

OpenStudy (wolf1728):

(2x+1) = log 8/log5 2x + 1 = 1.2920296742 2x = .2920296742 x = 0.1460148371

OpenStudy (wolf1728):

Testing the answer 5^(2*0.1460148371+1) = 8

OpenStudy (anonymous):

thank you!!!

OpenStudy (wolf1728):

no problem